# The Wavefront Set: Bounds for the Langlands parameter

### v2

@@ 5.2. Depth-zero representations
The canonical unramified wavefront sets of the depth-zero supercuspidal representations in the parametrization of [AX] are given in Table 2. In particular, Conjecture 1.1 holds for the depth-zero supercuspidals of $G_{2}$.
@@
The depth-zero supercuspidals are parameterized in loc. cit. as in the following Table 2. From the realization of the representations by compact induction, we can immediately compute the canonical unramified ${}^{K}\underline{\mathsf{WF}}(\pi)$ (in the sense of Okada [Ok]) and the geometric wavefront sets, via Theorem 3.1. Then one verifies by inspection the conjecture. The unipotent supercuspidal representations have already been treated in [CMBO2]. We recall that
$d^{\vee}(G_{2})=0,\quad d^{\vee}(0)=G_{2},\text{ and }d^{\vee}({\mathbb{O}}^{\vee})=G_{2}(a_{1}),$
if ${\mathbb{O}}^{\vee}=G_{2}(a_{1})$, $\widetilde{A}_{1}$, or $A_{1}$.
@@
the reflection representation of $W(G_{2})$. The corresponding nilpotent orbit via Springer’s correspondence is $G_{2}(a_{1}).$ This is the (geometric) Kawanaka wavefront set of $\mu$. Since the parahoric is the hyperspecial, it follows that the canonical unramified wavefront set of $\pi$ is indeed $(G_{2}(a_{1}),1)$.
@@
$G_{2}(a_{1})$ $\mu_{2}$ $SO_{4}$ singular $(G_{2}(a_{1}),(12))$ $G_{2}(a_{1})$ $\widetilde{A}_{1}$
$G_{2}(a_{1})$ $\mu_{3}$ $SL_{3}$ singular, $q\equiv 1(3)$ $(G_{2}(a_{1}),(123))$ $G_{2}(a_{1})$ $A_{1}$
$A_{1}$ $\mu_{3}$ $G_{2}$ singular from $SU(3)$ $(G_{2}(a_{1}),1)$ $G_{2}(a_{1})$ $G_{2}(a_{1})$
@@
$1$ $G_{2}$ nonsingular $(G_{2},1)$ $G_{2}$ $1$
$SL_{3}$ toral $(G_{2}(a_{1}),(123))$ $G_{2}$ $A_{1}$
$SO_{4}$ $(G_{2}(a_{1}),(12))$ $G_{2}$ $\widetilde{A}_{1}$
@@
To account for an example of an anti-tempered unipotent $E_{7}$-representation for which ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)<d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ (strict inequality), Okada proposed another, more precise form of the relation between the wavefront set of unipotent representations and the Langlands parameter:
$d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$ (5.1)
This holds for unipotent representations with real infinitesimal character [CMBO1], and there are no known counterexamples for unipotent representations. However, for more general representations, this equality can’t always hold as stated, for example, because of the case of nongeneric regular depth-zero supercuspidals, see section 3.1.
@@
${\mathbb{O}}^{\vee}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$ (5.2)
@@
More precisely, consider $\pi$ in the sixth row with ${\mathbb{O}}^{\vee}_{\pi}=A_{1}$ and ${}^{K}\underline{\mathsf{WF}}(\pi)=(G_{2}(a_{1}),1)$ and $d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))=G_{2}(a_{1})$. The intersection $\overline{d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))}\cap{\mathfrak{h}}^{\vee}$ is the principal orbit in $sl(3,{\mathbb{C}})$. However, the Langlands nilpotent orbit ${\mathbb{O}}^{\vee}=A_{1}$ comes from the subregular orbit in $sl(3,{\mathbb{C}})$.
@@
This calculation allows for an alternative verification of the conjecture in the case of depth-zero L-packets that contain a supercuspidal representation.
Proposition 5.5.
Conjecture 1.9 holds for all depth-zero discrete series L-packets of $G_{2}$ that contain a supercuspidal representation in the parametrization of [AX].
Proof.
The parametrization of the unipotent discrete series L-packets agrees with the Kazhdan-Lusztig parametrization [KL], hence the claim is already covered by [CMBO1] in the case of real infinitesimal character.
For the regular supercuspidal L-packets, the claim is immediate and has already been discussed in section 3.1
For the non-unipotent mixed L-packets, notice that Table 2 shows that the conjectural upper bound is attained by the supercuspidal in the L-packet. So it only remains to check that if $\pi$ is a non-supercuspidal representation in the L-packet, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}({\mathbb{O}}^{\vee})$. In all cases, $d^{\vee}({\mathbb{O}}^{\vee})=G_{2}(a_{1}).$ So the only way that the conjecture could fail is if ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=G_{2}$, which would make ${\mathsf{AZ}}(\pi)$ generic. But from the classification, we know that $\pi$ itself is generic, in fact $\pi$ is the generic composition factor of a generalized principal series representation that has two composition factors: $\pi$ and $\pi^{\prime}$. Then clearly ${\mathsf{AZ}}(\pi)=\pi^{\prime}$ which is not generic. This concludes the proof.
In fact, the previous proof can be adapted to conclude the following result (which is already known in more generality from [JLS, LoSa]).
Proposition 5.6.
Let $\pi$ be an irreducible unipotent representation of $G_{2}$. Then either $\pi$ is generic, or $\pi$ is trivial, or ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=G_{2}(a_{1})$.
Proof.
Suppose $\pi$ is not generic. Then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq G_{2}(a_{1})$. Consider the parahoric restriction $\mathsf{res}_{K_{0}}(\pi)$ of $\pi$ with respect to the maximal hyperspecial subgroup $K_{0}$. This is a unipotent representation of $G_{2}(\mathbb{F}_{q})$ and as such, its geometric Kawanaka wavefront set $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$ is special. The only special orbits in $G_{2}(\overline{\mathbb{F}}_{q})$ are the regular, subregular, and the trivial orbit. The only way $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$ is the trivial orbit is if $\mathsf{res}_{K_{0}}(\pi)$ is a trivial $G_{2}(\mathbb{F}_{q})$-representation. In particular, this means that $\pi$ is a $K_{0}$-spherical representation and in the equivalence of categories with the Iwahori-Hecke algebra, the corresponding Iwahori-Hecke algebra module only contains the trivial $K_{0}$-type. This implies that $\pi$ must be an unramified character, but since the group is $G_{2}$, it must be the trivial representation.
By [Ok, Theorem 5.4], ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ is greater than the lift to $G$ of $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$. So if $\pi$ is not trivial, it follows that ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\geq G_{2}(a_{1})$ and this completes the proof.

### v3

@@ 5.2. Unramified wavefront sets of depth-zero representations
The canonical unramified wavefront sets of the depth-zero supercuspidal representations in the parametrization of [AX] are given in Table 3.
@@
The depth-zero supercuspidals are parameterized in loc. cit. as in the following Table 3. From the realization of the representations by compact induction, we can immediately compute the canonical unramified $^{K}\underline{\mathsf{WF}}(\pi)$ (in the sense of Okada [Ok]) and the geometric wavefront sets, via Theorem 3.1. Then one verifies by inspection the conjecture. The unipotent supercuspidal representations have already been treated in [CMBO3]. We recall that
$d^{\vee}({\mathrm{G}_{2}})=1,\quad d^{\vee}(1)={\mathrm{G}_{2}},\text{ and }d^{\vee}({\mathbb{O}}^{\vee})={\mathrm{G}_{2}(a_{1})},$
if ${\mathbb{O}}^{\vee}={\mathrm{G}_{2}(a_{1})}$, $\widetilde{A}_{1}$}, or $A_{1}$}.
@@
the reflection representation of $W(G_{2})$. The corresponding nilpotent orbit via Springer’s correspondence is $\mathrm{G}_{2}(a_{1})$}. This is the (geometric) Kawanaka wavefront set of $\mu$. Since the parahoric is the hyperspecial, it follows that the canonical unramified wavefront set of $\pi$ is indeed $({\mathrm{G}_{2}(a_{1})},1)$.
@@
$\mathrm{G}_{2}(a_{1})$} $\mu_{2}$ $SO_{4}$ singular $({\mathrm{G}_{2}(a_{1})},(12))$ $\mathrm{G}_{2}(a_{1})$} $\widetilde{A}_{1}$
$\mathrm{G}_{2}(a_{1})$} $\mu_{3}$ $SL_{3}$ singular, $q\equiv 1(3)$ $({\mathrm{G}_{2}(a_{1})},(123))$ $\mathrm{G}_{2}(a_{1})$} $A_{1}$
$A_{1}$ $\mu_{3}$ $G_{2}$ singular from $SU(3)$ $({\mathrm{G}_{2}(a_{1})},1)$ $\mathrm{G}_{2}(a_{1})$} $\mathrm{G}_{2}(a_{1})$}
@@
11 $G_{2}$ nonsingular $({\mathrm{G}_{2}},1)$ $\mathrm{G}_{2}$} 11
$SL_{3}$ toral $({\mathrm{G}_{2}(a_{1})},(123))$ $\mathrm{G}_{2}$} $A_{1}$
$SO_{4}$ $({\mathrm{G}_{2}(a_{1})},(12))$ $\mathrm{G}_{2}$} $\widetilde{A}_{1}$
@@ 11 $G_{2}$ nonsingular $({\mathrm{G}_{2}},1)$ $\mathrm{G}_{2}$} 11
To account for an example of an anti-tempered unipotent $E_{7}$-representation for which $^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)<d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ (strict inequality), Okada proposed another, more precise form of the relation between the wavefront set of unipotent representations and the Langlands parameter [CMBO2, Conjecture 1.4.3]:
$d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{{\mathsf{AZ}}}.$ (5.1)
This holds for unipotent representations with real infinitesimal character [CMBO2], and there are no known counterexamples for unipotent representations. However, for more general representations, this equality can’t always hold as stated, for example, because of the case of nongeneric regular depth-zero supercuspidals, see section 3.1.
@@
${\mathbb{O}}^{\vee}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{AZ}}.$ (5.2)
@@
More precisely, consider $\pi$ in the sixth row of Table 3 with ${\mathbb{O}}^{\vee}_{\pi}=A_{1}$ and $^{K}\underline{\mathsf{WF}}(\pi)=({\mathrm{G}_{2}(a_{1})},1)$ and $d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))={\mathrm{G}_{2}(a_{1})}$. The intersection $\overline{d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))}\cap{\mathfrak{h}}^{\vee}$ is the principal orbit in $\mathfrak{sl}(3,{\mathbb{C}})$. However, the Langlands nilpotent orbit ${\mathbb{O}}^{\vee}=A_{1}$ comes from the subregular orbit in $\mathfrak{sl}(3,{\mathbb{C}})$.
@@
6. Acknowledgements
This paper builds on the previous work of the first-named author with Lucas Mason-Brown and Emile Okada. It is a pleasure to thank them for many useful conversations about the wavefront set. The second-named author thanks George Lusztig and Cheng-Chiang Tsai for insightful discussions. We also thank Baiying Liu and Freydoon Shahidi for sharing their work with us and for an interesting correspondence, and Hiraku Atobe and Alberto Minguez for clarifying the connection between Remark 1.11 and their AZ-algorithm [AM]. Special thanks are due to the referee for a very detailed and useful report that led to many improvements.
D.C. was partially supported by the EPSRC grant “New Horizons” EP/V046713/1, 1/2021-12/2023.
For the purpose of open access, the authors have applied a CC BY public copyright license to any author accepted manuscript arising from this submission.
On behalf of all authors, the corresponding author states that there is no conflict of interest.